# Sample Complexity of Linear Quadratic Regulator Without Initial Stability

### v2

@@ 7. Conclusion
First, according to Theorem 3.1, we select
(164) $N=\frac{1}{2}\cdot\frac{\log\big{(}\frac{2\|Q_{N}-P^{*}\|_{*}\cdot\kappa_{P^{*}}\cdot\|A_{K}^{*}\|\cdot\|B\|}{\epsilon\cdot\lambda_{\min}(R)}\big{)}}{\log\big{(}\frac{1}{\|A_{K}^{*}\|_{*}}\big{)}}+1,$
where $A_{K}^{*}:=A-BK^{*}$. This ensures that $K^{*}_{0}$ is stabilizing and $\|K^{*}_{0}-K^{*}\|\leq\epsilon/2$. Then, it remains to show that the output $\widetilde{K}_{0}$ satisfies $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$.
Recall that the RDE (20) is a backward iteration starting with $P^{*}_{N}=Q_{N}\geq 0$, and can also be represented as:
(165) $P^{*}_{t}$ $=(A-BK^{*}_{t})^{\top}P^{*}_{t+1}(A-BK^{*}_{t})+(K^{*}_{t})^{\top}RK^{*}_{t}+Q$
(166) $=A^{\top}P^{*}_{t+1}\big{(}A-BK^{*}_{t}\big{)}+Q+(K^{*}_{t})^{\top}(R+B^{\top}P^{*}_{t+1}B)K^{*}_{t}-(K^{*}_{t})^{\top}(B^{\top}P^{*}_{t+1}A)$
(167) =$\overset{\mathrm{(i)}}{=}A^{\top}P^{*}_{t+1}\big{(}A-BK^{*}_{t}\big{)}+Q,$
where (i) comes from $K^{*}_{t}=(R+B^{\top}P^{*}_{t+1}B)^{-1}(B^{\top}P^{*}_{t+1}A)$. Moreover, for clarity of proof, we denote the policy optimization error at time $t$ by:
$e_{t}:=\widetilde{K}_{t}-\widetilde{K}_{t}^{*}.$
We argue that $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$ can be achieved by carefully controlling $e_{t}$ for all $t$. At $t=0$, it holds that
$\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|+\|e_{0}\|,$
Hence, the error size $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|$ could be bounded by
Next, we require $\|e_{0}\|\leq\epsilon/4$ and $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\epsilon/4$ to fulfill $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$. We additionally require $\|P^{*}_{1}-\widetilde{P}_{1}\|\leq a$ to upper-bound the positive definite solutions of (18). Then, by (168), in order to fulfill $\|\widetilde{K}^{*}_{0}-K^{*}_{0}\|\leq\epsilon/4$, it suffices to require
(169) $\|P^{*}_{1}-\widetilde{P}_{1}\|\leq\min\bigg{\{}a,\frac{\epsilon}{4C_{1}}\bigg{\}}.$
Subsequently, we have
(170) $P^{*}_{1}-\widetilde{P}_{1}=(P^{*}_{1}-\widetilde{P}^{*}_{1})+(\widetilde{P}^{*}_{1}-\widetilde{P}_{1}).$
The first difference term on the RHS of (170) is
$P^{*}_{1}-\widetilde{P}^{*}_{1}$ $=A^{\top}P^{*}_{2}\big{(}A-BK^{*}_{1}\big{)}-A^{\top}\widetilde{P}_{2}\big{(}A-B\widetilde{K}^{*}_{1}\big{)}$
(171) $=A^{\top}(P_{2}^{*}-\widetilde{P}_{2})(A-BK^{*}_{1})+A^{\top}\widetilde{P}_{2}B(\widetilde{K}_{1}^{*}-K^{*}_{1}).$
(172) $=A^{\top}(P_{2}^{*}-\widetilde{P}_{2})(A-BK^{*}_{1})-A^{\top}\widetilde{P}_{2}B(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}(P^{*}_{2}-\widetilde{P}_{2})(A-BK^{*}_{1})$
(173) $=A^{\top}[I-\widetilde{P}_{2}B(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}](P_{2}^{*}-\widetilde{P}_{2})(A-BK^{*}_{1}),$
Moreover, the second term on the RHS of (170) is
$\widetilde{P}^{*}_{1}-\widetilde{P}_{1}$ $=(A-B\widetilde{K}^{*}_{1})^{\top}\widetilde{P}_{2}(A-B\widetilde{K}^{*}_{1})+(\widetilde{K}^{*}_{1})^{\top}R\widetilde{K}^{*}_{1}-(A-B\widetilde{K}_{1})^{\top}\widetilde{P}_{2}(A-B\widetilde{K}_{1})-(\widetilde{K}_{1})^{\top}R\widetilde{K}_{1}$
$=-(\widetilde{K}^{*}_{1})^{\top}B^{\top}\widetilde{P}_{2}A-A^{\top}\widetilde{P}_{2}B\widetilde{K}_{1}^{*}+(\widetilde{K}^{*}_{1})^{\top}(R+B^{\top}\widetilde{P}_{2}B)\widetilde{K}^{*}_{1}$
$\hskip 10.00002pt+\widetilde{K}_{1}^{\top}B^{\top}\widetilde{P}_{2}A+A^{\top}\widetilde{P}_{2}B\widetilde{K}_{1}-\widetilde{K}_{1}^{\top}(R+B^{\top}\widetilde{P}_{2}B)\widetilde{K}_{1}$
$=\big{[}(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}\widetilde{P}_{2}A-\widetilde{K}^{*}_{1}\big{]}^{\top}(R+B^{\top}\widetilde{P}_{2}B)\big{[}(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}\widetilde{P}_{2}A-\widetilde{K}^{*}_{1}\big{]}$
(174) $\hskip 10.00002pt-\big{[}(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}\widetilde{P}_{2}A-\widetilde{K}_{1}\big{]}^{\top}(R+B^{\top}\widetilde{P}_{2}B)\big{[}(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}\widetilde{P}_{2}A-\widetilde{K}_{1}\big{]}$
$=\big{[}\widetilde{K}^{*}_{1}-\widetilde{K}^{*}_{1}\big{]}^{\top}(R+B^{\top}\widetilde{P}_{2}B)\big{[}\widetilde{K}^{*}_{1}-\widetilde{K}^{*}_{1}\big{]}-\big{[}\widetilde{K}^{*}_{1}-\widetilde{K}_{1}\big{]}^{\top}(R+B^{\top}\widetilde{P}_{2}B)\big{[}\widetilde{K}^{*}_{1}-\widetilde{K}_{1}\big{]}$
(175) $=-e_{1}^{\top}(R+B^{\top}\widetilde{P}_{2}B)e_{1},$
where (174) follows from completion of squares. Thus, combining (170), (171), and (175) yields
$\hskip 13.00005pt\|P^{*}_{1}-\widetilde{P}_{1}\|$ $\leq\|P^{*}_{2}-\widetilde{P}_{2}\|\cdot\varphi\|A\|\|I-\widetilde{P}_{2}B(R+B^{\top}\widetilde{P}_{2}B)^{-1}B^{\top}\|+\|e_{1}\|^{2}\|R+B^{\top}\widetilde{P}_{2}B\|$
(176) $\leq\varphi\|A\|\left(1+\frac{\|P_{\max}+aI\|\|B\|^{2}}{\lambda_{\min}(R)}\right)\cdot\|P^{*}_{2}-\widetilde{P}_{2}\|+\|e_{1}\|^{2}\|R+B^{\top}\widetilde{P}_{2}B\|.$
Note that the difference between (176) with its counterpart in [20] is because the argument for
$\|(I+\widetilde{P}_{2}BR^{-1}B^{\top})^{-1}\|\leq 1$
does not hold, since $\widetilde{P}_{2}BR^{-1}B^{\top}$ is not necessarily positive semi-definite (a product of two symmetric psd matrices is not necessarily psd unless the product is a normal matrix). Now, we require
(177) $\|P^{*}_{2}-\widetilde{P}_{2}\|$ $\leq\min\bigg{\{}a,\frac{a}{C_{2}},\frac{\epsilon}{4C_{1}C_{2}}\cdot\bigg{\}}$
(178) $\|e_{1}\|$ $\leq\min\bigg{\{}\sqrt{\frac{a}{C_{3}}},\frac{1}{2}\sqrt{\frac{\epsilon}{C_{1}C_{3}}}\bigg{\}},$
Then, conditions (177) and (178) are sufficient for (169) (and thus for $\|\widetilde{K}_{0}-K^{*}_{0}\|\leq\epsilon/2$) to hold. Subsequently, we can propagate the required accuracies in (177) and (178) forward in time. Specifically, we iteratively apply the arguments in (176) (i.e., by plugging quantities with subscript $t$ into the LHS of (176) and plugging quantities with subscript $t+1$ into the RHS of (176)) to obtain the result that if at all $t\in\{2,\cdots,N-1\}$, we require
(179) $\|P^{*}_{t}-\widetilde{P}_{t}\|\leq\min\bigg{\{}a,\frac{a}{C_{2}^{t-1}},\frac{\epsilon}{4C_{1}C_{2}^{t-1}}\bigg{\}}$
$\|e_{t}\|\leq\min\bigg{\{}\sqrt{\frac{a}{C_{3}}},\sqrt{\frac{a}{C_{2}^{t-2}C_{3}}},\frac{1}{2}\sqrt{\frac{\epsilon}{C_{1}C_{2}^{t-2}C_{3}}}\bigg{\}},$
then (177) holds true and therefore (169) is satisfied.
We now compute the required accuracy for $e_{N-1}$. Note that $P^{*}_{N-1}=\widetilde{P}^{*}_{N-1}$ since no prior computational errors happened at $t=N$. By (176), the distance between $P^{*}_{N-1}$ and $\widetilde{P}_{N-1}$ can be bounded as
$\|P^{*}_{N-1}-\widetilde{P}_{N-1}\|=\|\widetilde{P}^{*}_{N-1}-\widetilde{P}_{N-1}\|\leq\|e_{N-1}\|^{2}\cdot C_{3}.$
To fulfill the requirement (179) for $t=N-1$, which is
$\|P^{*}_{N-1}-\widetilde{P}_{N-1}\|\leq\min\bigg{\{}a,\frac{a}{C_{2}^{N-2}},\frac{\epsilon}{4C_{1}C_{2}^{N-2}}\bigg{\}},$
it suffices to let
(180) $\|e_{N-1}\|\leq\min\bigg{\{}\sqrt{\frac{a}{C_{3}}},\sqrt{\frac{a}{C_{2}^{N-2}C_{3}}},\frac{1}{2}\sqrt{\frac{\epsilon}{C_{1}C_{2}^{N-2}C_{3}}}\bigg{\}}.$
Finally, we analyze the worst-case complexity of RHPG by computing, at the most stringent case, the required size of $\|e_{t}\|$. When $C_{2}\leq 1$, the most stringent dependence of $\|e_{t}\|$ on $\epsilon$ happens at $t=0$, which is of the order $\mathcal{O}(\epsilon)$, and the dependences on system parameters are $\mathcal{O}(1)$. We then analyze the case where $C_{2}>1$, where the requirement on $\|e_{0}\|$ is still $\mathcal{O}(\epsilon)$. Note that in this case, the requirement on $\|e_{N-1}\|$ is stricter than that on any other $\|e_{t}\|$ for any $t\in\{1,\cdots,N-1\}$ and by (180):
(181) $\|e_{N-1}\|\sim\mathcal{O}\Big{(}\sqrt{\frac{\epsilon}{C_{1}C_{2}^{N-2}C_{3}}}\Big{)}.$
Since we require $N$ to satisfy (164), the dependence of $\|e_{N-1}\|$ on $\epsilon$ in (181) becomes
$\|e_{N-1}\|\sim\mathcal{O}\left(\epsilon^{\frac{1}{2}+\frac{\log(C_{2})}{4\log{(1/\|A^{*}_{K}\|_{*}})}}\right)$
with additional polynomial dependences on system parameters since
$C_{2}^{N-2}$ $\approx C_{2}^{\frac{\log(1/\varepsilon)}{2\log(1/\|A^{*}_{K}\|_{*})}}$
$=(1/\varepsilon)^{\frac{\log(C_{2})}{2\log{\frac{1}{\|A^{*}_{K}\|_{*}}}}}.$
As a result, it suffices to require error bound for all $t$ to be
$\|e_{t}\|\sim\mathcal{O}\left(\min\left\{\varepsilon,\epsilon^{\frac{1}{2}+\frac{\log(C_{2})}{4\log{(1/\|A^{*}_{K}\|_{*}})}}\right\}\right)$
The difference between our requirement for the $C_{2}>1$ case with its counterpart in [20] is due to a calculation error in [20] which incorrectly neglects the impact of the exponent in $C_{2}^{N-2}$. Lastly, for $\widetilde{K}_{0}$ to be stabilizing, it suffices to select a sufficiently small $\epsilon$ such that the $\epsilon$-ball centered at the infinite-horizon LQR policy $K^{*}$ lies entirely in the set of stabilizing policies. A crude bound that satisfies this requirement is
$\epsilon<\frac{1-\|A-BK^{*}\|_{*}}{\|B\|}\Longrightarrow\|A-B\widetilde{K}_{0}\|_{*}<1.$
This completes the proof.

### v3

@@ 7. Conclusion
We follow [22, App. C] up to the recursion that controls the propagated error in the backward pass. First, select the horizon $N$ exactly as in Theorem 3.1, which ensures $\|K_{0}^{*}-K^{*}\|\leq\varepsilon/2$. Then, repeating their bookkeeping for the outer-loop error terms yields the familiar requirement on the last-stage inner-loop error
(160) $\|e_{N-1}\|\;\lesssim\;\mathcal{O}\left(\sqrt{\frac{\varepsilon}{\,C_{1}\,C_{2}^{\,N-2}\,C_{3}\,}}\right),$
which is precisely the counterpart of eq. (C.19) in [22].
Substituting the choice of $N$ from (31) into (160) gives
$C_{2}^{\,N-2}\approx C_{2}^{\,\frac{\log(1/\varepsilon)}{2\log(1/\|A^{*}_{K}\|_{*})}}=(1/\varepsilon)^{\frac{\log C_{2}}{2\log(1/\|A^{*}_{K}\|_{*})}},$
and hence
(161) $\|e_{N-1}\|\sim\mathcal{O}\left(\varepsilon^{\frac{1}{2}+\frac{\log C_{2}}{4\log(1/\|A^{*}_{K}\|_{*})}}\right).$
In [22], this exponent is (incorrectly) simplified to $3/4$, i.e., as if $\frac{\log C_{2}}{4\log(1/\|A^{*}_{K}\|_{*})}=\frac{1}{4}$ held universally; however, this does not hold, since this ratio is system-dependent in general. The same substitution appears explicitly in[22, (C.4) and (C.19)]. Indeed, this error propagates to their final sample-complexity. Since the inner-loop PG (two-point estimation) needs $\widetilde{\mathcal{O}}(\|e_{h}\|^{-2})$ steps to reach the required accuracy at stage $h$, the overall iteration budget must scale as
$\widetilde{\mathcal{O}}\left(\max\big\{\|e_{0}\|^{-2},\|e_{N-1}\|^{-2}\big\}\right)$
$\hskip 50.00008pt=\widetilde{\mathcal{O}}\!\left(\max\!\left\{\varepsilon^{-2},\,\varepsilon^{-1-\frac{\log C_{2}}{2\log(1/\|A^{*}_{K}\|_{*})}}\right\}\right),$
rather than $\widetilde{\mathcal{O}}(\varepsilon^{-2})$ obtained by substituting $\|e_{N-1}\|=\mathcal{O}(\varepsilon^{3/4})$.
Our fix. Our outer-loop analysis avoids the exponential factor $C_{2}^{N-2}$ entirely by exploiting that the Riccati operator $\mathcal{R}(\cdot)$ is nonexpansive (contractive) in the Riemannian distance $\delta(\cdot,\cdot)$ under Assumption 2.1 (Lemma 2.1). This yields stage-wise requirements that do not blow up with $N$, recovering the $\widetilde{\mathcal{O}}(\varepsilon^{-2})$ total rate with a one-point model-free gradient estimator.